40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  CTN Lab elog  Not logged in ELOG logo
Entry  Tue Jul 2 11:21:24 2013, Erica, DailyProgress, fiber optic, mode matching snip_of_beam_profiling_6_27.JPGP1020518.JPG
    Reply  Wed Jul 3 09:46:48 2013, Erica, DailyProgress, fiber optic, mode matching 130703-095332.jpg
       Reply  Thu Jul 11 01:40:19 2013, Erica, DailyProgress, fiber optic, mode matching P1020690.JPGP1020692.JPG
          Reply  Fri Jul 12 00:51:39 2013, Erica, Notes, fiber optic, mode matching P1020705.JPGP1020703.JPG7_11_13_beam_profile.PNG7_11_13_beam_from_fixed_lens.PNG
             Reply  Tue Jul 16 10:27:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 
                Reply  Tue Jul 16 11:02:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 130805-120628.jpg
                   Reply  Wed Jul 17 11:58:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 7_16_13_determining_polarization_axis.PNGTEK0000.PNG
                      Reply  Wed Jul 17 18:29:37 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                         Reply  Sat Jul 20 23:39:09 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                            Reply  Tue Jul 23 10:29:12 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam shielded_10s_TEK0005.PNGunshielded_10s_TEK0027.PNG
                               Reply  Tue Jul 23 11:43:00 2013, Evan, DailyProgress, Electronics Equipment, Zach's op-amp recommendations 
                               Reply  Fri Jul 26 11:07:44 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table P1030056.JPGP1030057.JPGP1030059.JPGP1030056.JPG
                                  Reply  Fri Jul 26 11:22:45 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                     Reply  Fri Jul 26 11:45:28 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                        Reply  Mon Jul 29 13:30:26 2013, Erica, DailyProgress, BEAT, adding resistor, capacitor, and sockets w_o_capacitor_7_26_13.JPGcapacitor_7_26_13.JPGP1030066.JPGP1030070.JPG
                                           Reply  Tue Jul 30 12:23:41 2013, Erica, DailyProgress, BEAT, addressed noise problem, DC signal 
                                              Reply  Thu Aug 1 00:25:02 2013, Erica, DailyProgress, fiber optic, insulating foam 731_7graph.PNG731_11graph.PNGP1030318.JPG
                                                 Reply  Fri Aug 2 12:38:41 2013, tara, DailyProgress, fiber optic, insulating foam chmbrtmp.JPGfringe.JPG
                                                    Reply  Tue Aug 6 12:31:24 2013, Erica, DailyProgress, fiber optic, insulating foam P1030403.JPG
                                                       Reply  Wed Aug 7 13:27:00 2013, Erica, DailyProgress, fiber optic, more foam, calculations for convertion from voltage to frequency P1030409.JPGP1030417.JPGP1030420.JPG
                                                          Reply  Wed Aug 7 18:53:20 2013, Erica, DailyProgress, fiber optic, noise budget, Matlab notes. noise_budget.PNGP1030427.JPG
                                                             Reply  Tue Aug 13 12:39:36 2013, Erica, DailyProgress, fiber optic, rewrote calculations for converting from voltage noise to frequency noise P1030458.JPGP1030459.JPGP1030460.JPG
                                                       Reply  Tue Aug 13 12:28:20 2013, Erica, Notes, Electronics Equipment, new AD 590 circuit and its noise calculation, beam alignment P1030461.JPGP1030454.JPGP1030456.JPGP1030457.JPGAD590circuit2.PNG
                                                          Reply  Fri Aug 16 10:35:44 2013, Erica, Notes, Electronics Equipment, AD590 msrmt, conversion to freq noise 
                   Reply  Tue Jul 30 16:19:32 2013, Erica, Notes, fiber optic, setup for beam recombination new_setup.PNGbeam_recombination2.PNG
                   Reply  Mon Aug 5 12:21:00 2013, Erica, Notes, fiber optic, calculation: phase change for fiber 130805-121449.jpg
    Reply  Tue Jul 30 15:48:47 2013, Erica, DailyProgress, fiber optic, noise for GYRO 
Message ID: 1222     Entry time: Tue Jul 2 11:21:24 2013     Reply to this: 1224   1266
Author: Erica 
Type: DailyProgress 
Category: fiber optic 
Subject: mode matching 
Sorry I type up my progress, but keep on forgetting to submit.

June 27, 2013

Talked to Evan to better understand GYRO. The goal of GYRO is to be able to differentiate between tilt and actual noise. There is a laser that is split and goes in opposite directions around a cavity which has mirrors on opposite corners and beam splitters on the other 2 corners; the laser is locked to the cavity. This setup uses the Sagnac effect. We are sending the CTN laser over because it has a lower noise, due to the shorter cavity, and will serve as reference for the noise level for the GYRO setup. I emailed Zach to ask about more specifics on the GYRO noise because he has a better idea but he has not yet replied.

I used a WinCam and a program called DataRay to look at the Gaussian profile of the beam at various differences from the output of the fiber(see setup here: http://nodus.ligo.caltech.edu:8080/PSL_Lab/1213). I took the radii from the two profiles at each distance, averaged them, and plotted that as a function of distance from the lens in the collimator and photodiode in the WinCam. This is to find the beam waist and its location because the current setup did not result in a beam waist at a location we predicted. See attachment 1.

I fit the equation 20 from Kogelnik and Li, 1966 to the graph and obtained a waist of 35um, at z = -0.0237 m. To check, I used equation 22, theta = lambda/(pi*w0) and the slope of the graph, to solve for w0. This ended up agreeing with the 35um.



June 28, 2013


Read papers Tara gave. Looked around for ones on related topics.

Took the new waist from yesterday's data and found a new setup. BPS program doesn't seem to work. After closing and reopening JamMT, 4 lenses show up on the diagram despite the fact I only chose 2 lenses, so I'm not sure what was going on. I ended up just doing trial and error in the program, until I got a setup that fit my parameters.

Setup is in attachment 2. The lens with f= 85.8mm is the lens that is already on the table. I added lenses with f=74.9mm and 143.2mm to mode match.

The beam waist and its location seems to match what was calculated.

I played around with different locations and got a maximum output of 48um. Unfortunately, I didn't look at the power further up the beam so I don't know what percentage. The power output at various locations seem rather strange because the data didn't converge to a single maximum, despite having made the most adjustments I could at each point.



July 1, 2013

Tried to find out more about self-homodyne detection and noise that the fiber optic contributes. Tried to find Eric Q but can't seem to ever find him so I emailed him to talk about crackle lab.

Went to Alan's talk today on the sources of gravitational waves. They include compact binaries, core collapses, spinning neutron stars, and cosmic gravitational wave background.

With 2 mW power, I was able to obtain 100uW from setup seen in attachment 1. I set up a half-wave plate, beam splitter, and mirrors to be able to recombine the beam tomorrow.
Attachment 1: snip_of_beam_profiling_6_27.JPG  78 kB  | Hide | Hide all
snip_of_beam_profiling_6_27.JPG
Attachment 2: P1020518.JPG  924 kB  | Hide | Hide all
P1020518.JPG
ELOG V3.1.3-